OpenStudy (anonymous):

sin^2 theta - cos^2 theta = 1

OpenStudy (solomonzelman):

the identity is \(\large\color{black}{ \rm sin^2 θ + cos^2θ = 1 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf sin^2 θ - cos^2θ = 1 }\) \(\large\color{black}{ \bf sin^2 θ - (1-sin^2θ) = 1 }\) \(\large\color{black}{ \bf sin^2 θ - 1+sin^2θ = 1 }\) \(\large\color{black}{ \bf sin^2 θ+sin^2θ = 1+1 }\) \(\large\color{black}{ \bf 2sin^2θ = 2 }\) \(\large\color{black}{ \bf sin^2θ = 1 }\) \(\large\color{black}{ \bf sinθ = ± 1 }\) \(\large\color{black}{ \bf sin^{-1}1 =θ ~~~~~~and~~~~~~\large\color{black}{ \bf sin^{-1}1 =θ }}\)

OpenStudy (solomonzelman):

I mean ^(-1)1, and ^(-1) -1

zepdrix (zepdrix):

I like solomon's method :) Another approach is to use the Cosine Double Angle Formula. Just something else to try :D\[\Large\rm \cos^2x-\sin^2x=-1,\]\[\Large\rm \cos2x=-1,\]\[\Large\rm 2x=\pi+2k \pi,\]Solving for x gives,\[\Large\rm x=\frac{\pi}{2}+k \pi\]